# Fuzzy Classification Aggregation

### v4

@@ 1 Introduction
The study of the problem of individuals classifying objects can be traced back to Kasher and Rubinstein, (1997). In their paper, they consider a finite society that has to determine which one of its subsets of members consists of exactly those individuals that can be deemed to be part of a group named $J$. They propose different sets of axioms characterizing three aggregators: the liberal one, where each individual decides by herself whether she belongs or not to the group; the oligarchic one, where a subset of individuals defines who belongs to the group; and the dictatorial, where a unique individual decides who is a $J$. Kasher and Rubinstein, (1997) gave rise to what can be regarded as a subdomain within social choice theory, namely, the study of the Group Identification Problem (see, among others, Samet and Schmeidler,, 2003; Miller,, 2008; Fioravanti and Tohmé,, 2021, for more details).
@@
Our work considers the problem where $m$ objects have to be classified into $p$ categories, $m\geq p\geq 2$, and the classifications, both by the individuals and the rule, indicate degrees of membership of the objects to each of the categories. Two possible scenarios where this setting may arise, are the following. A national government has to allocate equal funds to all regions, to be used for different regional ministries (security, education, and health), with the central government providing some conditions on how the funds are distributed. It is natural to assume in this context, that the government sets a minimum amount of money that should be spent in each of the ministries, at least at a national level. Members of the government should decide on how the funds might be allocated for security, education, and health, ensuring that each category, across the country, receives at least the amount provided by the government to each region. As an illustration, one region could be tasked with allocating half of the funds to security, a quarter to education, and a quarter to health. Another scenario involves assigning individuals to various mandatory tasks with equal time requirements, where proportions dictate how each person’s time is allocated. For instance, if tasks include laundry, lawn mowing, and grocery shopping, Bob might spend half of his time on laundry, a quarter on mowing the lawn, and the remaining quarter on grocery shopping. [^3] These examples illustrate why the notion of “degree of membership” can be considered natural to use, instead of considering probabilities. For instance, in the tasks allocation example, it is better if we know that Bob will spend with certainty half of his time doing the laundry, rather than knowing that there is a $50\%$ chance of Bob leaving all the clothes dirty.
@@
We avoid both impossibility results, the crisp and the fuzzy ones, by considering a different, but also very natural, fuzzy surjectivity condition. This condition, which in the crisp setting is extremely demanding, becomes a requirement that in the fuzzy setting allows the existence of a family of aggregators satisfying certain particular properties. Thus, it is demanding enough to find a characterization result, but not so much to not be able to find one. Building on some previous results in functional analysis by Aczél and Wagner, (1980); Aczél et al., (1984), and Wagner, (1982), we show that the only optimal aggregators that are independent and that satisfy a weak version of Unanimity, are the Weighted Arithmetic Means.
Considering different settings for already established problems is a common technique used in social choice to avoid impossibility results. The result by Gibbard, (1973) and Satterthwaite, (1975), that states in the deterministic setting that the only rule immune to manipulations is the dictatorial one, is avoided when the outcomes of the rule are lotteries over the set of alternatives (Gibbard,, 1977). We consider that our paper is related to Maniquet and Mongin, (2016)’s paper in an analogous way as Dutta, (1987) and Billot, (1991) results are related to the Arrow,’s (1951) impossibility theorem. In their work, they consider a fuzzy setting and adapt Arrow’s theorems to it to obtain possibility results. Dutta, (1987) characterizes the oligarchic aggregators, while Billot, (1991) does the same for the leximin aggregator. The use of a fuzzy setting, that allows many interpretations of axioms that are used in crisp settings, proves advantageous for circumventing impossibility results.
The plan of the paper is as follows. Section 2 presents the basic notions and axioms that we use, while we present the results in Section 3. Finally, Section 4 contains some concluding remarks.
@@ 2 Basic Notions and Axioms
Let $N=\{1,\ldots,n\}$ be a finite set of individuals and let $X=\{x_{1},\ldots,x_{m}\}$ be $m$ objects that need to be classified into the $p$ categories of a set $P$, with $p\geq 2$.
The individuals classify each object according to a partial degree of membership to each category. In the crisp setting, introduced by Maniquet and Mongin, (2016), classifications are surjective mappings $c:X\rightarrow P$, that is, every category must have at least one object classified into. In the fuzzy setting, a fuzzy classification is a mapping $c:X\rightarrow[0,1]^{P}$ such that $\sum_{j=1}^{m}c(x_{j})_{t}\geq 1$ for each $t\in\{1,\ldots,p\}$ and $\sum_{t=1}^{p}c(x)_{t}=1$ for all $x\in X$. The former condition is the fuzzy counterpart of the surjectivity of the classification function, while the latter is the fuzzy counterpart of the assumption that every object $x$ must be assigned to exactly one category by each voter. It is easy to see that these conditions imply that there must be at least as many objects as categories, thus $m\geq p$.
@@
$\alpha_{\bm{w}}(\bm{c})(x)=w_{1}c_{1}(x)+\cdots+w_{n}c_{n}(x)$
We say that the set of weights is degenerate if there is an $i\in N$ such that $w_{i}=1$. If we think of the FCAF as a group of experts classifying objects, this FCAF can be appropriate for situations where the individuals differ in expertise, with more experienced voters having higher weights. [^4] Weighted aggregators have been previously studied in the social choice literature, in particular in the setting where monetary transfers are allowed (see Roberts,, 1979; Mishra and Sen,, 2012). See also Stone, (1961) and Dietrich and List, (2016) for an use of this aggregator in opinion pooling. In particular, if $w_{i}=\frac{1}{n}$ for all $i\in N$, we call this FCAF the Arithmetic Mean.
@@
An FCAF is optimal if for all $\bm{c}\in\mathcal{C}^{N}$ such that $\sum_{j=1}^{m}c_{i}(x_{j})_{t}=h$ for all $i\in N$, it is the case that $\sum_{j=1}^{m}\alpha(\bm{c})(x_{j})_{t}=h$.
This axiom is useful in a setting where categories need to fulfil certain requirements and resources should not be wasted. For example, in the case where a set of managers has to classify a group of workers as members of different departments within a company, in order to make them operative. If all the managers consider that department A requires only two workers to be operative, then an optimal aggregator only classifies two workers into that department. The following axiom states that the fuzzy social classification of an object in two different fuzzy classification profiles does not change if the classification regarding that object is the same in both profiles for every individual.
@@
An FCAF is independent if given $\bm{c,c^{\prime}}\in\mathcal{C}^{N}$ and $x\in X$ such that $c_{i}(x)=c^{\prime}_{i}(x)$ for all $i\in N$, it is the case that $\alpha(\bm{c})(x)=\alpha(\bm{c^{\prime}})(x)$.
An independent FCAF $\alpha$ can be seen as a collection of mappings $(\alpha_{x})_{x\in X}$, such that $\alpha_{x}:\mathcal{C}^{N}|_{x}\rightarrow\mathcal{C}|_{x}$ and $\alpha_{x}(\bm{c}^{x})=\alpha(\bm{c})(x)$ (Cailloux et al.,, 2024). We call these mappings Elementary FCAFs. A stronger version of Independence requires that if two objects are classified equally in a profile, then the fuzzy social classification must be the same. [^5] This property is called Strong Label Neutrality in Wagner, (1982).
@@
An FCAF is symmetric if for all $x,y\in X$ and all $\bm{c}\in\mathcal{C}^{N}$ such that $c_{i}(x)=c^{\prime}_{i}(y)$ for all $i\in N$, it is the case that $\alpha(\bm{c})(x)=\alpha(\bm{c})(y)$.
Wagner, (1982) shows that an optimal and symmetric FCAF is also independent, and thus has the same elementary FCAF for every object. The next property states that if there is an object that is unanimously classified by the individuals, then the FCAF has to classify that object accordingly.
@@
An FCAF is unanimous if for all $\bm{c}\in\mathcal{C}^{N}$ where there is an $x\in X$ and a category $t\in P$ such that $c_{1}(x)_{t}=\cdots=c_{n}(x)_{t}=r$, it is the case that $\alpha(\bm{c})(x)_{t}=r$.
@@
An FCAF is zero unanimous if for all $\bm{c}\in\mathcal{C}^{N}$ where there is an $x\in X$ and a category $t\in P$ such that $c_{1}(x)_{t}=\cdots=c_{n}(x)_{t}=0$, it is the case that $\alpha(\bm{c})(x)_{t}=0$.
@@
An FCAF satisfies Fuzzy Consensus if for every object $x\in X$ and for every category $t\in\{1,\ldots,p\}$, it is the case that $\alpha(\mathbf{c})(x)_{t}\in[\min_{i\in N}c_{i}(x)_{t},\max_{i\in N}c_{i}(x)_{t}]$.
It is easy to see that Fuzzy Consensus implies Unanimity, and Unanimity implies Zero Unanimity. These three axioms can be interpreted as the different ‘degrees’ of consensus that we might require an aggregator to satisfy. The next axiom states that there must not exist an individual that imposes her classification.
@@
An FCAF is non-dictatorial if there is no individual $i\in N$ such that for all $x\in X$ and all $\bm{c}\in\mathcal{C}^{N}$, it is the case that $\alpha(\mathbf{c})(x)=c_{i}(x)$.
A stronger version of Non-Dictatorship states that the names of the individuals are not important for the aggregation of classifications.
@@ 3 Results
if, and only if, there exists a set of weights $(w_{1},\ldots,w_{n})$ such that for $j=1,\ldots,m$ and any $z\in[0,1]^{N}$, we have that
@@
That a Weighted Arithmetic Mean with a non-degenerate set of weights satisfies the four axioms is straightforward to see. For the only if part, we use the result by Aczél et al., (1984). Let $\alpha$ be an optimal FCAF that satisfies Independence and Zero Unanimity. Recall that an independent FCAF can be seen as a collection of elementary FCAFs. So we can consider a family of mappings $\{\alpha_{x_{j}}:\mathcal{C}^{N}|_{x_{j}}\rightarrow\mathcal{C}|_{x_{j}}\}_{j=1}^{m}$. By optimality, we obtain that if $\sum_{j=1}^{m}c_{i}(x_{j})_{t}=h$ for all $i\in N$, it is the case that $\sum_{j=1}^{m}\alpha_{x_{j}}(\bm{c}^{x_{j}})_{t}=h$. Then we have that for each $t\in P$ there is a family of mappings $\{\alpha_{x_{j}}|_{t}:\mathcal{C}^{N}|_{x_{j}\land t}\rightarrow[0,1]\}_{j=1}^{m}$ satisfying condition (1). The function $\alpha_{x_{j}}|_{t}$ maps the column $t$ in $\bm{c}^{x_{j}}$ and outputs the degree of classification of object $x_{j}$ into the category $t$. Due to optimality, regardless of the values that are not in column $t$, we can be sure that condition (1) is satisfied by $\alpha_{x_{j}}|_{t}$ and that the degree of classification of the object $x_{j}$ into the category $t$ depends on a weighted aggregator. Thus we have that for each category $t$ there is a set of weights $(w^{t}_{1},\ldots,w^{t}_{n})$ such that for $j=1,\ldots,m$, it is the case that
@@
What is left to show is that for every $i\in N$, the weights are the same for every category, meaning that $w_{i}^{t}=w_{i}^{t^{\prime}}$ for all $t,t^{\prime}\in P$. To avoid excessive notation, we show it for the case of $|N|=2$, and $m=p=3$. It is easy to extend it for the general case. Consider the example where $c_{1}(x_{1})=(1,0,0)$, $c_{1}(x_{2})=(0,1,0)$, $c_{1}(x_{3})=(0,0,1)$, $c_{2}(x_{1})=(0,1,0)$, $c_{2}(x_{2})=(0,0,1)$, and $c_{2}(x_{3})=(1,0,0)$. Thus we obtain $\alpha_{x_{1}}(c_{1}(x_{1}),c_{2}(x_{1}))=(w_{1}^{1},w_{2}^{2},0)$, $\alpha_{x_{2}}(c_{1}(x_{2}),c_{2}(x_{2}))=(0,w_{1}^{2},w_{2}^{3})$, and $\alpha_{x_{3}}(c_{1}(x_{3}),c_{2}(x_{3}))=(w_{2}^{1},0,w_{1}^{3})$, where every three dimensional vector adds up to one. Thus we obtain the following system of equations:
$\begin{cases}w_{1}^{1}+w_{2}^{1}=1&(1)\\ w_{1}^{2}+w_{2}^{2}=1&(2)\\ w_{1}^{3}+w_{2}^{3}=1&(3)\\ w_{1}^{1}+w_{2}^{2}=1&(4)\\ w_{1}^{2}+w_{2}^{3}=1&(5)\\ w_{2}^{1}+w_{1}^{3}=1&(6)\end{cases}$
From $(1)$ and $(4)$ we obtain that $11-w_{2}^{1}+w_{2}^{2}=$ 1, thus $w_{2}^{1}=w_{2}^{2}$, and from $(2)$ and $(5)$ we obtain that $11-w_{2}^{2}+w_{2}^{3}=$ 1, thus $w_{2}^{2}=w_{2}^{3}$. It is a similar procedure for the weights of agent $1$. So the FCAF associates to every individual $i\in N$ a weight $w_{i}$ such that $w_{i}\geq 0$ and that $\sum_{i=1}^{n}w_{i}=1$, and as it is non-dictatorial, we have that there is no $i\in N$ such that $w_{i}=1$. Thus the FCAF is a Weighted Arithmetic Mean with a non-degenerate set of weights. ∎
@@
Let $m\geq 3$ and $m\geq p\geq 2$. An optimal Fuzzy Classification Aggregation Function satisfies Independence, Zero Unanimity, and Anonymity if, and only if, it is an Arithmetic Mean.
@@
Thus, both impossibility results by Maniquet and Mongin, (2016) and Alcantud et al., (2019) are avoided in our setting, even if we use their same ‘consensual’ axioms.
@@
A more general result, valid for $m\geq p\geq 2$ can be attained if we slightly change the surjectivity conditions. For this general case, we consider fuzzy classifications $c^{\ast}:X\rightarrow\mathbb{R}^{P}$, and the surjectivity requirements are such that for all $x\in X$, and for a given $s\in\mathbb{R}$, it is the case that $\sum_{t=1}^{p}c^{\ast}(x)_{t}=s$, and that $\sum_{j=1}^{m}c^{\ast}(x_{j})_{t}\geq s$ if $s\geq 0$ for all $t\in P$, or $\sum_{j=1}^{m}c^{\ast}(x_{j})_{t}<s$ if $s<0$ for all $t\in P$. We can think of this setting as $n$ operators assigning $s$ hours of use of $m$ machines into $p$ different tasks that need $s$ hours to be finished, or $s$ euros that $m$ persons have to spend in $p$ different projects that need at least $s$ euros to be done. We use $\mathcal{C}^{\ast}$ to denote this set of classifications. Thus, the FCAF is a mapping $\alpha:{\mathcal{C}^{\ast}}^{N}\rightarrow\mathcal{C}^{\ast}$ such that $\alpha(\bm{c}^{\ast})(x_{j})$ indicates, for example, the different hours assigned to the different tasks of the machine $x_{j}$.
We have that an optimal FCAF satisfies the $k$-allocation property for any $k=m\geq 2$ (Aczél and Wagner,, 1980), that is, if for any $t\in P$ such that $\sum_{j=1}^{m}c^{\ast}(x_{j})_{t}=s$ for all $i\in N$, it is the case that $\sum_{j=1}^{m}\alpha(\bm{c}^{\ast})(x_{j})_{t}=s$. The $k$-allocation property allows us to characterize the optimal and symmetric FCAFs that satisfy Fuzzy Consensus for any value of $m$ such that $m\geq p\geq 2$.
@@
Aczél and Wagner, (1980) show that a mapping $\bm{c}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ satisfies the k-allocation property for $k=2,3$ and is bounded if, and only if, it is a Weighted Arithmetic Mean. In our setting, symmetry implies that for a given category $t\in P$, we have the same elementary FCAF for every object. By optimality, the k-allocation property is satisfied by the elementary FCAFs. Fuzzy Consensus implies they are bounded and thus they are Weighted Arithmetic Means, and finally, Non-Dictatorship implies that the set of weights is non-degenerate. By a similar proof to the one used in Theorem 1, it is easy to see that the elementary FCAFs are the same for every category, concluding our proof. ∎

### v5

@@ 1 Introduction
The study of the problem of individuals classifying objects can be traced back to Kasher and Rubinstein, (1997). In their paper, they consider a finite society that has to determine which one of its subsets of members consists of exactly those individuals that can be deemed to be part of a group named $J$. Kasher and Rubinstein, (1997) gave rise to what can be regarded as a subdomain within social choice theory, namely, the study of the Group Identification Problem (see, among others, Samet and Schmeidler,, 2003; Miller,, 2008; Fioravanti and Tohmé,, 2021, for more details).
@@
Our work considers the problem where $m$ objects have to be classified into $p$ categories, $m\geq p\geq 2$, and the classifications, both by the individuals and the rule, indicate degrees of membership of the objects to each of the categories. Two possible scenarios where this setting may arise, are the following. A national government has to allocate equal funds to all regions, to be used for different regional ministries (security, education, and health), with the central government providing some conditions on how the funds are distributed. It is natural to assume in this context, that the government sets a minimum amount of money that should be spent in each of the ministries, at least at a national level. Members of the government should decide on how the funds might be allocated for security, education, and health, ensuring that each category, across the country, receives at least the amount provided by the government to each region. As an illustration, one region could be tasked with allocating half of the funds to security, a quarter to education, and a quarter to health. Another scenario involves assigning individuals to various mandatory tasks with equal time requirements, where proportions dictate how each person’s time is allocated. For example, if tasks include laundry, lawn mowing, and grocery shopping, Bob, one of the individuals assigned to do the tasks, might spend half of his time on laundry, a quarter on mowing the lawn, and the remaining quarter on grocery shopping. [^3] These examples illustrate why the notion of “degree of membership” can be considered natural to use, instead of considering probabilities. For instance, in the tasks allocation example, it is better if we know that Bob will spend with certainty half of his time doing the laundry, rather than knowing that there is a $50\%$ chance of Bob leaving all the clothes dirty.
@@
We avoid both impossibility results, the crisp and the fuzzy ones, by considering a different, but also very natural, fuzzy surjectivity condition. This condition, which in the crisp setting is extremely demanding, becomes a requirement that in the fuzzy setting allows the existence of a family of aggregators satisfying certain particular properties. Thus, it is demanding enough to find a characterization result, but not so much to not be able to find one. Building on some previous results in functional analysis by Aczél and Wagner, (1980); Aczél et al., (1984), and Wagner, (1982), we show that the only optimal aggregators that are independent and that satisfy a weak version of Unanimity, are the Weighted Arithmetic Means (WAMs). These rules aggregate individual inputs by computing a weighted average of their classifications, with weights reflecting the relative influence of each individual.
The Weighted Arithmetic Means are particularly significant for their theoretical and practical implications. Theoretically, they provide a framework that is both mathematically elegant and normatively appealing. By resolving prior impossibility results, they demonstrate that meaningful aggregation is achievable in the fuzzy setting under natural surjectivity conditions. Practically, their versatility and interpretability make them well-suited to a wide range of real-world decision-making problems.
In economic contexts, the Weighted Arithmetic Means offer a powerful tool for balancing efficiency and equity. For example, they can model scenarios where agents differ in expertise, authority, or stake, such as resource allocation in government planning, voting systems in elections, or task assignment in organisational settings. The weights assigned to individuals can reflect various factors, such as their expertise in evaluating specific categories, the resources they contribute, or their perceived reliability.
Consider the previously introduced example where a national government has to allocate equal funds to all regions, but now the members of the government are also representative of the different regions. A WAM-based approach allows to aggregate regional preferences into a collective allocation while respecting the relative importance or influence of each region. Similarly, in collaborative decision-making within organisations or project teams, WAMs facilitate the fair integration of expert opinions by assigning higher weights to individuals with specialized knowledge, ensuring that critical insights are not overlooked.
Furthermore, WAMs embody a natural interpretation of fairness: individuals with equal expertise or standing are treated equally, while those with greater expertise are appropriately weighted to reflect their contributions. This fairness property is particularly valuable in settings where the credibility or importance of agents varies, ensuring that outcomes remain justifiable and transparent.
The significance of these results lies not only in their theoretical robustness but also in their practical applicability to diverse fields, including economics, political science, and artificial intelligence, where fuzzy and nuanced classifications are increasingly common. By providing a clear axiomatic foundation, the Weighted Arithmetic Means offer a bridge between theory and application, enabling decision-makers to navigate complex classification problems in a principled and effective manner.
The plan of the paper is as follows. Section 2 presents the basic notions and axioms that we use, while we present the results in Section 3. The Subsection 3.1 we include examples that show the independence of the axioms of our main result, Theorem 1. Finally, Section 4 contains some concluding remarks.
@@ 2 Basic Notions and Axioms
Let $N=\{1,\ldots,n\}$ be a finite set of individuals and let $X=\{x_{1},\ldots,x_{m}\}$ be a set of $m$ objects that need to be classified into the $p$ categories of a set $P$, with $p\geq 2$.
The individuals classify each object according to a partial degree of membership to each category. In the crisp setting, introduced by Maniquet and Mongin, (2016), classifications are surjective mappings $c:X\rightarrow P$, that is, every category must have at least one object classified into. In the fuzzy setting, a fuzzy classification is a mapping $c:X\rightarrow[0,1]^{P}$ such that $\sum_{j=1}^{m}c(x_{j})_{t}\geq 1$ for all $t\in P$ and $\sum_{t=1}^{p}c(x)_{t}=1$ for all $x\in X$. The former condition is the fuzzy counterpart of the surjectivity of the classification function, while the latter is the fuzzy counterpart of the assumption that every object $x$ must be assigned to exactly one category by each voter. It is easy to see that these conditions imply that there must be at least as many objects as categories, thus $m\geq p$.
@@
$\alpha_{\bm{w}}(\bm{c})(x)=w_{1}c_{1}(x)+\cdots+w_{n}c_{n}(x).$
We say that the set of weights is degenerate if there is an $i\in N$ such that $w_{i}=1$. If we think of the FCAF as a group of experts classifying objects, the WAM can be appropriate for situations where the individuals differ in expertise, with more experienced voters having higher weights. [^4] Weighted aggregators have been previously studied in the social choice literature, in particular in the setting where monetary transfers are allowed (see Roberts,, 1979; Mishra and Sen,, 2012). See also Stone, (1961) and Dietrich and List, (2016) for an use of this aggregator in opinion pooling. In particular, if $w_{i}=\frac{1}{n}$ for all $i\in N$, we call this FCAF the Arithmetic Mean.
@@
An FCAF is optimal if for all $\bm{c}\in\mathcal{C}^{N}$ and all $t\in P$ such that $\sum_{j=1}^{m}c_{i}(x_{j})_{t}=h$ for all $i\in N$, it is the case that $\sum_{j=1}^{m}\alpha(\bm{c})(x_{j})_{t}=h$.
This axiom is useful in a setting where categories need to fulfil certain requirements and resources should not be wasted. For example, in the case where a set of managers has to classify a group of workers as members of different departments within a company, in order to make them operative. If all the managers consider that department A requires only two workers to be operative, then an optimal aggregator only classifies two workers into that department.
The following axiom states that the fuzzy social classification of an object in two different fuzzy classification profiles does not change if the classification regarding that object is the same in both profiles for every individual.
@@
An FCAF is independent if for all $\bm{c},\bm{c^{\prime}}\in\mathcal{C}^{N}$ and all $x\in X$ such that $c_{i}(x)=c^{\prime}_{i}(x)$ for all $i\in N$, it is the case that $\alpha(\bm{c})(x)=\alpha(\bm{c^{\prime}})(x)$.
This property is desirable in settings such as budget allocations, where the classification of one item (e.g., education spending) should not depend on unrelated factors (e.g., infrastructure spending). Independence is a relatively weak requirement, as it asserts that the values of certain variables should be disregarded, in a setting where constraints are already imposed on the sums of these variables. We can note that an independent FCAF $\alpha$ can be seen as a collection of mappings $(\alpha_{x})_{x\in X}$, such that $\alpha_{x}:\mathcal{C}^{N}|_{x}\rightarrow\mathcal{C}|_{x}$ and $\alpha_{x}(\bm{c}^{x})=\alpha(\bm{c})(x)$ (Cailloux et al.,, 2024). We call these mappings Elementary FCAFs.
A stronger version of Independence requires that if two different objects are classified equally in two different fuzzy classification profiles, then the fuzzy social classification must be the same for both objects. [^5] This property is called Strong Label Neutrality in Wagner, (1982).
@@
An FCAF is symmetric if for all $x,y\in X$ and all $\bm{c},\bm{c^{\prime}}\in\mathcal{C}^{N}$ such that $c_{i}(x)=c^{\prime}_{i}(y)$ for all $i\in N$, it is the case that $\alpha(\bm{c})(x)=\alpha(\bm{c^{\prime}})(y)$.
Besides making the classifications of the rest of objects irrelevant when classifying, for example, an object $x_{j}$, a symmetric FCAF also ignores the names of the objects when classifying them. In economic contexts, this can be seen as a fairness requirement, ensuring that identical cases are treated equally. For example, in labour market classifications, two identical workers should receive the same classification, regardless of arbitrary labels. Wagner, (1982) shows that an optimal and symmetric FCAF has the same elementary FCAF for every object.
The next property states that if there is an object that is unanimously classified by the individuals, then the FCAF has to classify that object accordingly.
@@
An FCAF is unanimous if for all $\bm{c}\in\mathcal{C}^{N}$, all $x\in X$, all $t\in P$, and all $r\in[0,1]$ such that $c_{1}(x)_{t}=\cdots=c_{n}(x)_{t}=r$, it is the case that $\alpha(\bm{c})(x)_{t}=r$.
This property aligns with fundamental democratic principles in economic policy; if all stakeholders agree on a classification, it should be implemented.
@@
An FCAF is zero unanimous if for all $\bm{c}\in\mathcal{C}^{N}$, all $x\in X$, and all $t\in P$ such that $c_{1}(x)_{t}=\cdots=c_{n}(x)_{t}=0$, it is the case that $\alpha(\bm{c})(x)_{t}=0$.
This property can also be seen as a rationality constraint; if no member of the government wants to allocate funds for security in a specific region, then that region should not receive funds for that ministry.
@@
An FCAF satisfies Fuzzy Consensus if for all $\bm{c}\in\mathcal{C}^{N}$, all $x\in X$, and all $t\in P$, it is the case that $\alpha(\mathbf{c})(x)_{t}\in[\min_{i\in N}c_{i}(x)_{t},\max_{i\in N}c_{i}(x)_{t}]$.
With an FCAF that satisfies Fuzzy Consensus, we can be confident that an object’s classification remains close to the classifications assigned by the individuals, preventing extreme or arbitrary outcomes.
It is easy to see that Fuzzy Consensus implies Unanimity, and Unanimity implies Zero Unanimity. These three axioms can be interpreted as the different ‘degrees’ of consensus that we might require an aggregator to satisfy.
The next axiom states that there must not exist an individual that imposes her classification. It is useful to prevent autocratic outcomes and ensure the participation of all the individuals.
@@
An FCAF is non-dictatorial if there is no individual $i\in N$ such that $\alpha(\mathbf{c})=c_{i}$ for all $\bm{c}\in\mathcal{C}^{N}$.
A stronger version of Non-Dictatorship states that the names of the individuals are not important for the aggregation of classifications. This property ensures fairness and prevents undue influence based on identity rather than the merit of classifications.
@@ 3 Results
if, and only if, there exists a set of weights $(w_{1},\ldots,w_{n})$ such that for each $j=1,\ldots,m$ and any $z\in[0,1]^{N}$, we have that
@@
[…]
@@
What is left to show is that for every $i\in N$, the weights are the same for every category, meaning that $w_{i}^{t}=w_{i}^{t^{\prime}}$ for all $t,t^{\prime}\in P$.
Fix an $i^{\star}\in N$ and consider a set of classification problems $\{\bm{c}^{t}\}_{1\leq t\leq p}$ such that for each $\bm{c}^{t}$, it is the case that $c^{t}_{j}(x_{1})=(1,0,\ldots,0)$ for all $j\neq i^{\star}$ and $c^{t}_{i^{\star}}(x_{1})=(0,\ldots,0,\underset{t-\text{th}}{1},0,\ldots,0)$. Thus, as for each classification problem $\bm{c}^{t}$ we have that $\alpha(\bm{c}^{t})(x_{1})=(\sum_{j\neq i^{\star}}w_{j}^{1},\ldots,0,\underset{t-\text{th}}{w_{i^{\star}}^{t}},0,\ldots,0)$, we obtain the constraint
$\sum_{j\neq i^{\star}}w_{j}^{1}+w_{i^{\star}}^{t}=1\hskip 56.9055pt(t)$
In consequence, for any pair of classification problems $\bm{c}^{t},\bm{c}^{t^{\prime}}$, we can substract the constrains $(t)$ and $(t^{\prime})$ and obtain that $w_{i^{\star}}^{t}=w_{i^{\star}}^{t^{\prime}}$. This conclusion holds for all $t,t^{\prime}\in P$, and for each $i\in N$.
So the FCAF associates to every individual $i\in N$ a weight $w_{i}$ such that $w_{i}\geq 0$ and that $\sum_{i=1}^{n}w_{i}=1$, and as it is non-dictatorial, we have that there is no $i\in N$ such that $w_{i}=1$. Thus the FCAF is a Weighted Arithmetic Mean with a non-degenerate set of weights. ∎
@@
Let $m\geq 3$ and $m\geq p\geq 2$. An optimal Fuzzy Classification Aggregation Function satisfies Independence, Zero Unanimity, and Anonymity if, and only if, it is the Arithmetic Mean.
@@
The impossibility results established by Maniquet and Mongin, (2016) and Alcantud et al., (2019) are successfully circumvented in our framework, even while preserving their original ‘consensual’ axioms. This breakthrough is primarily due to our refined and less restrictive definition of a surjective category, which serves as the foundation for our results. By weakening the surjectivity conditions proposed in these earlier works in a natural and intuitive manner, we demonstrate that properties like Independence and Unanimity, which are often regarded as overly demanding, become more attainable in the context of fuzzy classifications.
@@
A more general result, valid for $m\geq p\geq 2$ can be attained with some slightly changes in the model. For this general case, let $A\subseteq X$, $|A|=a\geq 2$, and consider fuzzy classifications $c^{\ast}_{A}:A\rightarrow\mathbb{R}^{P}$, where the surjectivity requirements are such that for all $x\in A$, and for a given $s\in\mathbb{R}$, it is the case that $\sum_{t=1}^{p}c_{A}^{\ast}(x)_{t}=s$, and that $\sum_{j=1}^{a}c_{A}^{\ast}(x_{j})_{t}\geq s$ if $s\geq 0$ for all $t\in P$, or $\sum_{j=1}^{a}c_{A}^{\ast}(x_{j})_{t}<s$ if $s<0$ for all $t\in P$. We can think of this setting as $n$ operators assigning $s$ hours of use of at most $m$ machines into $p$ different tasks that need $s$ hours to be finished, or $s$ euros that at most $m$ persons have to spend in $p$ different projects that need at least $s$ euros to be done. This setting can be relevant when not always all the machines or all the persons are available, but we need the decision process to be consistent across the different possibilities. For each $A\subseteq X$ we use $\mathcal{C}_{A}^{\ast}$ to denote this set of classifications. We define an FCAF $\alpha$ as a set of FCAFs over subsets of $X$, that is, $\alpha=\{\alpha_{A}\}_{A\subseteq X}$, where $\alpha_{A}:{\mathcal{C}^{\ast}_{A}}^{N}\rightarrow\mathcal{C}^{\ast}_{A}$. We say that an FCAF is optimal if $\alpha_{A}$ is optimal for all $A\subseteq X$. We similarly define a symmetric, fuzzy consensual and non-dictatorial FCAF.
We have that an optimal FCAF satisfies the $k$-allocation property for any $k=a\geq 2$ (Aczél and Wagner,, 1980), that is, if for any $t\in P$ such that $\sum_{j=1}^{a}c^{\ast}_{A}(x_{j})_{t}=s$ for all $i\in N$, it is the case that $\sum_{j=1}^{m}\alpha(\bm{c}_{A}^{\ast})(x_{j})_{t}=s$. The $k$-allocation property allows us to characterize the optimal and symmetric FCAFs that satisfy Fuzzy Consensus for all values of $m$ such that $m\geq p\geq 2$.
@@
Aczél and Wagner, (1980) show that a mapping $\bm{c}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ satisfies the k-allocation property for $k=2,3$ and is bounded if, and only if, it is a Weighted Arithmetic Mean. In our setting, symmetry implies that for a given category $t\in P$ and a given $A\subseteq X$, $\alpha_{A}$ can be seen as a set of elementary FCAFs $\{\alpha_{x}\}_{x\in A}$, where $\alpha_{x}=\alpha_{y}$ for all $x,y\in A$. By optimality, the k-allocation property is satisfied by $\alpha_{A}$ for all $A\subseteq X$, in particular for the cases where $|A|=2$ or $|A|=3$. Fuzzy Consensus implies they are bounded and thus, by Aczél and Wagner, (1980), the mapping $\alpha_{A}$ is a Weighted Arithmetic Mean for all $A\subseteq X$. Finally, Non-Dictatorship implies that the set of weights is non-degenerate. By a similar proof to the one used in Theorem 1, it is easy to see that the mappings $\alpha_{A}$ are the same for every category, concluding our proof. ∎
3.1 Independence of the Axioms in Theorem 1
We conclude this section by showing that the axioms in our main result are independent. For each of the axioms in the statement of Theorem 1, we exhibit an example of an FCAF, different from a Weighted Arithmetic Mean, that satisfies all the axioms but one.
  • All but Optimality: let $m=3$ and $p=2$. The FCAF $\alpha$ such that for all $\bm{c}\in\mathcal{C}^{N}$, for any $x\in X$ such that $c_{1}(x)=(1,0)$ or $c_{1}(x)=(0,1)$, then $\alpha(\bm{c})(x)=c_{1}(x)$; otherwise it is $\alpha(\bm{c})(x)=(\frac{1}{2},\frac{1}{2})$. This FCAF is almost always constant, except when the individual $1$ proposes extreme classifications.
  • All but Independence: the FCAF $\alpha$ such that if $\bm{c}\in\mathcal{C}^{N}$ is such that for all $i\in N$, $c_{i}(x_{1})=e_{i}$ (the canonical vector), then $\alpha(\bm{c})=c_{1}$; otherwise it is $\alpha(\bm{c})=c_{2}$.
  • All but Zero Unanimity: the FCAF $\alpha$ such that for all $\bm{c}\in\mathcal{C}^{N}$, $\alpha(\bm{c})(x_{1})=c_{1}(x_{2})$, $\alpha(\bm{c})(x_{2})=c_{1}(x_{1})$, and $\alpha(\bm{c})(x_{j})=c_{1}(x_{j})$ for $j\neq 1,2$. This FCAF is essentially a dictatorship (Cailloux et al.,, 2024).
  • All but Non-Dictatorship: the FCAF such that for all $\bm{c}\in\mathcal{C}^{N}$, it is the case that $\alpha(\mathbf{c})=c_{i}$ for all $\bm{c}\in\mathcal{C}^{N}$, for a fixed $i\in N$.